40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 19 18:27:16 2015, ericq, Update, ASC, QPD ASC saga continues.  oplevPlant.pdfloopDesignComparison.pdfJan16_newLoops_offOn.png
    Reply  Tue Jan 20 02:39:49 2015, ericq, Update, ASC, QPD ASC saga continues.  
    Reply  Tue Feb 10 18:37:17 2015, ericq, Update, ASC, QPD ASC ready to go Feb10_loops_offOn.pngFeb10_newLoops_offOn.xml.zip
       Reply  Tue Feb 24 18:54:26 2015, ericq, Update, ASC, Single arm QPD ASC stability trend.png
Message ID: 10997     Entry time: Tue Feb 10 18:37:17 2015     In reply to: 10920     Reply to this: 11061
Author: ericq 
Type: Update 
Category: ASC 
Subject: QPD ASC ready to go 

I've remeasured the QPD ASC sensing coefficients, and figured out what I did weird with the actuation coefficients. I've rearranged the controller filters to be able to turn on the boost in a triggered way, and written Up/Down scripts that I've tested numerous times, and Jenne has used as well; they are exposed on the ASC screen. 

All four loops (2 arms * pit,yaw), have their gains set for 8Hz UGF, and have 60 degrees of phase margin. The loop shape is the same as the previous ELOG post. Here is the current on/off performance. The PDH signals (not shown, but in attached xml) show no extra noise, and the low frequency RIN goes down a bit, whic is good. The oplevs error signals are a bit noisier, but I suppose that's unavoidable. The Y-arm performs a bit better than the X-arm. 

The up/down scripts don't touch the filters' trigger settings at all, just handles switching the input and output and clearing history. FM1 contains the boost which is intended to have a longer trigger delay than the filters themselves.

Attachment 1: Feb10_loops_offOn.png  81 kB  | Hide | Hide all
Feb10_loops_offOn.png
Attachment 2: Feb10_newLoops_offOn.xml.zip  1.969 MB
ELOG V3.1.3-